OpenStudy (anonymous):

Help pleaaaseeee.. 1) Find the balance in the account $2000 principle earning 7% compounded semi-annually after 25 years. 2)A tractor costs $14,340 and depreciates in value by 15% per year. How much will the tractor be worth after 3 years?

OpenStudy (anonymous):

i forgot how to do 1 but i can help on 2

OpenStudy (anonymous):

\[21000\times \left(1+\frac{.07}{2}\right)^{50}\] for the first one that, and a calculator

OpenStudy (anonymous):

okay since it depreciates by 15 every year for 3 years you would multiply .15 by 3 ( i put 15% in decimal form then it's more easier to calculate)

OpenStudy (anonymous):

.15 times 3 is .45

OpenStudy (anonymous):

now you would do 14,340 times .45 which is $6,453

OpenStudy (anonymous):

got it??

OpenStudy (anonymous):

yepp no problem!! Happy to help

OpenStudy (anonymous):

unfortunately it is wrong

OpenStudy (anonymous):

way wrong sorry

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

to decrease a number by \(15\%\) you multiply it by \(100\%-15\%=85\%=.85\)

OpenStudy (anonymous):

to do it three times in a row, once for each year, you multiply it by \(.85^3\)

OpenStudy (anonymous):

the answer you are looking for is found by computing \[14,340\times (.85)^3\]

OpenStudy (anonymous):

can u please help me and explain it to me more tomorrow.. @satellite73